# Quantum Buffer Design Using Petri Nets

### v1

@@ 3.2 Multiple Input/Single Output Quantum Buffer
A MISO (Multiple Input, Single Output) quantum buffer is a type of quantum buffer that has many entries and only one exit point for data $q$-tokens. The MISO quantum buffer allows pairing any of its $k$ data inputs (places) with its only output place to transfer a $q$-token from that input to the output. Figure 7 depicts the QPN model of MISO quantum buffer with capacity $m,$ $k+1$ transitions $T_{1},T_{2},...T_{k+1}$, $k+4$ places $P_{I1},P_{I2},...P_{Ik},P_{DA},P_{A},P_{A1},P_{O}$, $r_{j}$ data $q$-tokens $d_{1,1},d_{1,2},...,d_{j,r_{j}}$ in $P_{Ij},1\leq j\leq k,$ and $m$ ancillary $q$-tokens $z_{1},...,z_{m}$. Each $z_{i}$ is specified as an address $q$-token that is used to select one of the transitions $T_{1},T_{2},...T_{k}$ to fire in order to transfer a data $q$-token in one of the input places $P_{I1},P_{I2},...P_{Ik}$ to output place $P_{O},$ via place $P_{DA}.$ Each of $z_{1},...,z_{m}$ causes one transition to fire and the MISO buffer halts its operation when all $m$ address $q$-tokens are utilized. The labels on the directed arcs $x_{1},x_{2},\cdots,x_{3k+3}$ denote the variables that represent data $q$-tokens or pairs of data and address $q$-tokens on the incoming and outgoing arcs of transitions. The qubits on the outgoing arcs from $P_{A}$ denote the address $q$-tokens that serve as selectors of transitions to which the outgoing arcs connect place $P_{A}.$ We note that only one transition fires at a time so that $P_{DA}$ receives two $q$-tokens (one data and one address $q$-token) at any point in time. The $q$-tokens are assigned to places $P_{I1},P_{I2},...P_{Ik},$ and $P_{A}$ are initialized before the buffer starts operating as follows:
Figure 7: A multiple-input-single-output (MISO) quantum buffer.
@@
When the input transition $T_{i},$ fires $1\leq i\leq k,$ it forwards the current data and ancillary $q$-token, say $d_{i,y}$ and $z_{y}$ to place $P_{DA}$ i.e., $f_{1}$ becomes $d_{i,y},z_{y}$. When the output transition $T_{k+1}$ fires, it forwards the current data and ancillary $q$-token to $P_{O}$ and $P_{A1},$ respectively. The behavior of the MISO quantum buffer of capacity $m=2$ with two input places $P_{I1}$ and $P_{I2}$ is illustrated in Figure 8. The transitions are fired according to the address values of the ancillary qubits and the number of $q$-tokens in the input places. For example, transition $T_{1}$ can fire twice along a given sequence of transitions since $m=2$ and place $P_{I1}$ contains two $q$-tokens. On the other hand, transition $T_{2}$ can only fire once as place $P_{I2}$ has only one $q$-token. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. Some transitions cannot be fired due to $q$-token constraints. To process $m$ data $q$-tokens, $2m$ steps (transitions) are required, the whole process is reversible, and can be repeated again.
Figure 8: The flow of $q$-tokens in a quantum MISO buffer, $m=2.$
@@ 3.3 Single Input/Multiple Output Quantum Buffer
A SIMO (Single Input, Multiple Output) quantum buffer is a type of quantum buffer that has only one entry and many exit points for data $q$-tokens. It should be pointed out that the data $q$-tokens in input places are not copied to multiple output places all at once, but each data $q$-token is forwarded to a selected output, one at a time. $P_{DA}$ during each transition can hold up to $m$ pairs of data and ancillary $q$-tokens. Figure 9 depicts the QPN model of SIMO quantum buffer of capacity $m,$ composed of $k$ transitions $T_{1},T_{2},...T_{k}$, $k+3$ places $P_{I},P_{A},P_{A1},P_{O1},P_{O2},...{P}_{Ok}$, $n$ data $q$-tokens $d_{1},d_{2},...,d_{n},$ and $m$ ancillary $q$-tokens $z_{1},...,z_{m},$ each of which can be assigned a quantum value from $|0\!>$ to $|k-1\!\!>$. The labels on the directed arcs $x_{1},x_{2},\cdots,x_{3k}$ denote the variables that represent $q$-tokens on the incoming and outgoing arcs of transitions.
@@
The initial $q$-tokens are assigned to places $P_{I}$ and $P_{A}$ as in SISO quantum buffer. When transition $T_{j},1\leq j\leq k$ fires, it forwards the current data and address $q$-tokens to $P_{Oj}$ and $P_{A1},$ respectively. The behavior of the SIMO buffer of capacity $m=3$ with two output places $P_{O1}$ and $P_{O2}$ is illustrated in Figure 10.
Figure 10: The flow of $q$-tokens in a quantum SIMO buffer, $m=2.$
The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. We note that at the end of possible transitions, the four data $q$-tokens may be distributed to $P_{O1}$ and $P_{O2}$ in one of four ways: (3,0), (2,1), (1,2), (0,3) as the bottom row in Figure 10 shows. Also, the value of $m$ limits the number of transitions to 3 in this example, preventing transitions with output patterns (4,0) (3,1), (2,2), (1,3), (0,4). The diagram shows only the feasible sequences of transitions, and to process $m$ data $q$-tokens, $m$ steps (transitions) are needed, where the whole process is reversible, and can be repeated again.
@@ 3.4 Multiple Input Multiple Output Quantum Buffer
A MIMO (Multiple Input, Multiple Output) quantum buffer is a type of quantum buffer that has multiple entry and multiple exit points for data $q$-tokens. Two sets of ancillary $q$-tokens are used to select the input and output places as the source and destination of a pair of transitions. The ancillary $q$-tokens $w_{1},w_{2},\cdots,w_{m}$ are used to select an input place, and ancillary $q$-tokens $z_{1},z_{2},\cdots,z_{m}$ are used to select an output place. Figure 11 depicts the QPN model of MIMO quantum buffer of capacity $m,$ composed of $s$ transitions $T_{1},T_{2},T_{3},...,T_{s}$, $k$ input places $P_{I1},P_{I2},...,P_{Ik},s-k$ output places $P_{O1},P_{O2},...,P_{s-k}$, two ancillary places $P_{A},P_{A1},$ one joint data/ancillary place $P_{DA},$ $n=r_{1}+r_{2}+\cdots+r_{k}$ data $q$-tokens $d_{1,1},d_{1,2},...,d_{k,rk},$ and $2m$ ancillary $q$-tokens $w_{1},w_{2},\cdots,w_{m},z_{1},...,z_{m}$.
@@
The labels on the directed arcs $x_{1},x_{2},\cdots,x_{4s}$ denote the variables that represent $q$-tokens on the incoming and outgoing arcs of transitions. The data $q$-tokens are assigned to places $P_{I1},P_{I2},...P_{Ik},$ and $P_{A}$ as in MISO quantum buffer, and not repeated here. When the input transition $T_{i}$ fires, $1\leq i\leq k,$ it forwards the current data and ancillary $q$-token, say $d_{i,y}$ and $z_{y}$ to place $P_{DA}$ i.e., $f_{1}$ becomes $d_{i,y},z_{y}$. When the output transition $T_{j}$ fires, $k+1\leq j\leq s,$ it forwards an ancillary $q$-token $z_{y}$ and an ancillary $q$-token $w_{y}$ to $P_{A2},$ and data $q$-token $d_{i,y},$ to the corresponding output place $P_{oj}$. The behavior of the MIMO quantum buffer of capacity $m=2$ with two input and two output places $P_{I1},P_{I2},P_{O1}$ and $P_{O2}$ is illustrated in Figure 12. Input places $P_{I1}$ and $P_{I2}$ are initialized to two and one $q$-tokens, respectively.
Figure 12: The flow of $q$-tokens in a quantum MIMO buffer, $m=2.$
The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. We note that at the end of possible transitions, the two $q$-tokens may be distributed from input places $P_{I1},P_{I2},$ to output places $P_{O1}$ and $P_{O2}$ in one of six ways: (0,1,2,0), (0,1,1,1), (0,1,0,2), (1,0,2,0), (1,0,1,1), (1,0,0,2) as the diagram in Figure 12 shows. Note that some transition sequences result in identical patterns of $q$-tokens in output places and are counted once. For example, both $T_{1},T_{3},T_{1},T_{4}$ and $T_{1},T_{4},T_{1},T_{3}$ result in the placement of one token in each of $P_{o1}$ and $P_{o1}.$ It takes $2m$ transitions, to process $m$ data $q$-tokens. Overall, the whole process is reversible, and can be repeated again.
@@ 3.5 Priority Quantum Buffer
A priority quantum buffer is a type of quantum buffer that prioritizes the processing of some $q$-tokens over others. Both data and ancillary $q$-tokens are divided into two groups: high priority and low priority. The high priority ancillary $q$-tokens are employed to manage the prioritization of data $q$-tokens that reside in $P_{I2}$ over the data $q$-tokens in $P_{I1}$ during processing. It should be pointed out that the input transitions $T_{1},T_{2}$ can run independently. Figure 13 depicts the QPN model of priority quantum buffer with capacity $m,$ four transitions $T_{1},T_{2},T_{3},T_{4}$, eight places $P_{I1},P_{I2},P_{DA1},P_{A},P_{DA2},P_{A1},P_{A2},P_{O}$, $r$ low priority data $q$-tokens, $d_{1},d_{2},...,d_{r}$, $n-r$ high priority data $q$-tokens $d_{r+1},d_{r+2},...,d_{n}$, $r+1\leq j\leq n,$ $m$ low priority ancillary $q$-tokens (qubits), $w_{1},w_{2},...,w_{m}$ and $m$ high priority ancillary $q$-tokens (qubits), $z_{1},z_{2},...,z_{m}$. The labels on the directed arcs $x_{1},x_{2},\cdots,x_{13}$ denote the variables that represent $q$-tokens on the incoming and outgoing arcs of transitions. The directed arc $x_{9}$ also known as inhibitor arc, prevents the firing of output transition $T_{3}$ if there is any $q$-token present in $P_{DA2}$. The $q$-tokens are assigned to places $P_{I1},P_{I2},$ and $P_{A1}$ and are initialized before the buffer starts operating as follows:
@@
Figure 14: The flow of $q$-tokens in a priority quantum buffer, $m=3.$
When the input transition, say $T_{i}$ fires, $1\leq i\leq 2,$ it forwards the current data and ancillary $q$-token, say $d_{y}$ and $z_{y}$ (or $w_{y}$) to place $P_{DAi}$ i.e., $f_{i}$ becomes either $d_{y},z_{y}$ (or $d_{y},w_{y}$). When the output transition, say $T_{j}$ fires, $3\leq j\leq 4,$ it forwards the corresponding ancillary $q$-token $z_{y}$ or $w_{y}$ to $P_{A2}$ and data $q$-token $d_{y}$ to $P_{O}$. However, the data $q$-tokens in place $P_{DA2}$ are prioritized over all tokens in place $P_{DA1}$ at the output place $P_{O}$ because transition $T_{3}$ cannot fire, due to the constraint of the inhibitor arc $x_{9}$. The behavior of the priority quantum buffer of capacity $m=4$ is illustrated in Figure 14. Input places $P_{I1}$ and $P_{I2}$ are initialized to one and two $q$-tokens, respectively. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. We note that some transition sequences inhibit the firing of transition $T_{3}$. For example, when $T_{2}$ fires, it inhibit the firing of transition $T_{3}$ unless $T_{4}$ fired. It takes $2m$ transitions, to process $m$ data $q$-tokens. Overall, the whole process is reversible, and can be repeated again.
4 Quantum Buffer Examples
In this section we test the QPN buffer model by implementing the Quantum SR flip-flop and SISO, SIMO, MISO, MIMO and priority quantum buffer designs on the IBM composer [29] and Jupyter [30] using Python and Qiskit [31].

### v2

@@ 4.2 Multiple Input/Single Output Quantum Buffer
A MISO (Multiple Input, Single Output) quantum buffer is a type of quantum buffer that has many entries and only one exit point for data $q$-tokens. The MISO quantum buffer allows pairing any of its $k$ data inputs (places) with its only output place to transfer a $q$-token from that input to the output. Figure 8 depicts the QPN model of MISO quantum buffer with capacity $m,$ $k$ transitions $T_{1},T_{2},...T_{k}$, $k+3$ places $P_{I1},P_{I2},...P_{Ik},P_{A},P_{A1},P_{O}$, $r_{j}$ data $q$-tokens $d_{1,1},d_{1,2},...,d_{j,r_{j}}$ in $P_{Ij},1\leq j\leq k,$ and $m$ ancillary $q$-tokens $z_{1},...,z_{m}$. Each $z_{i}$ is specified as an address $q$-token that is used to select one of the transitions $T_{1},T_{2},...T_{k}$ to fire in order to transfer a data $q$-token in one of the input places $P_{I1},P_{I2},...P_{Ik}$ to output place $P_{O}.$ This is accomplished [^1] Conditional transitions are permitted in quantum petri nets[25]. by a condition placed on each link that compares the current $z_{i},1\leq i\leq m$ token against the qubits that are marked on the arcs between $P_{A}$ and $T_{i},1\leq i\leq k.$ Each of $z_{1},...,z_{m}$ causes one transition to fire and the MISO buffer halts its operation when all $m$ address $q$-tokens are consumed. The labels on the directed arcs $x_{1},x_{2},\cdots,x_{4k+4}$ denote the variables that represent data $q$-tokens or pairs of data and address $q$-tokens on the incoming and outgoing arcs of transitions. The $q$-tokens are assigned to places $P_{I1},P_{I2},...P_{Ik},$ and $P_{A}$ are initialized before the buffer starts operating as follows:
@@
Figure 8: A multiple-input-single-output (MISO) quantum buffer.
The behavior of the MISO quantum buffer of capacity $m=2$ with two input places $P_{I1}$ and $P_{I2}$ is illustrated in Figure 9. The transitions are fired according to the address values of the ancillary qubits and the number of $q$-tokens in the input places. For example, transition $T_{1}$ can fire twice along a given sequence of transitions since $m=2$ and place $P_{I1}$ contains two $q$-tokens. On the other hand, transition $T_{2}$ can only fire once as place $P_{I2}$ has only one $q$-token. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. Some transitions cannot be fired due to $q$-token constraints. To process $m$ data $q$-tokens, $m$ steps (transitions) are required, the whole process is reversible, and can be repeated again.
Figure 9: The flow of $q$-tokens in a quantum MISO buffer, $k=2,r_{1}=2,r_{2}=1,m=3.$
@@ 4.3 Single Input/Multiple Output Quantum Buffer
A SIMO (Single Input, Multiple Output) quantum buffer is a type of quantum buffer that has only one entry and many exit points for data $q$-tokens. Figure 10 depicts the QPN model of SIMO quantum buffer of capacity $m,$ composed of $k$ transitions $T_{1},T_{2},...T_{k}$, $k+3$ places $P_{I},P_{A},P_{A1},P_{O1},P_{O2},...{P}_{Ok}$, $n$ data $q$-tokens $d_{1},d_{2},...,d_{n},$ and $m$ ancillary $q$-tokens $z_{1},...,z_{m},$ each of which can be assigned a quantum value from $|0\!>$ to $|k-1\!\!>$. The labels on the directed arcs $x_{1},x_{2},\cdots,x_{4k}$ denote the variables that represent $q$-tokens on the incoming and outgoing arcs of transitions. The initial $q$-tokens are assigned to places $P_{I}$ and $P_{A}$ as in SISO quantum buffer. When transition $T_{j},1\leq j\leq k$ fires, it forwards the current data and address $q$-tokens to $P_{Oj}$ and $P_{A1},$ respectively. The behavior of the SIMO buffer of capacity $m=3$ with two output places $P_{O1}$ and $P_{O2}$ is illustrated in Figure 11. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. We note that at the end of possible transitions, the four data $q$-tokens may be distributed to $P_{O1}$ and $P_{O2}$ in one of four ways: (3,0), (2,1), (1,2), (0,3) as the bottom row in Figure 11 shows. Also, the value of $m$ limits the number of transitions to 3 in this example, preventing transitions with output patterns (4,0) (3,1), (2,2), (1,3), (0,4). The diagram shows only the feasible sequences of transitions, and to process $m$ data $q$-tokens, $m$ steps (transitions) are needed, where the whole process is reversible, and can be repeated again.
@@
Figure 11: The flow of $q$-tokens in a quantum SIMO buffer, $k=2,m=3,n\geq 3.$
@@ 4.4 Multiple Input Multiple Output Quantum Buffer
A MIMO (Multiple Input, Multiple Output) quantum buffer is a type of quantum buffer that has multiple entry and multiple exit points for data $q$-tokens. Two sets of ancillary $q$-tokens are used to select the input and output places as the source and destination of a pair of transitions. The ancillary $q$-tokens $w_{1},w_{2},\cdots,w_{m}$ are used to select a $q$-token from an input place, and ancillary $q$-tokens $z_{1},z_{2},\cdots,z_{m}$ are used to select an output place to which the selected $q$-token is transferred. Figure 12 depicts the QPN model of MIMO quantum buffer of capacity $m,$ composed of $s$ transitions $T_{1},T_{2},T_{3},...,T_{s}$, $k$ input places $P_{I1},P_{I2},...,P_{Ik},s-k$ output places $P_{O1},P_{O2},...,P_{O(s-k)},$ two ancillary places $P_{A},P_{A1},$ one joint data/ancillary place $P_{DA},$ $n=r_{1}+r_{2}+\cdots+r_{k}$ data $q$-tokens $d_{1,1},d_{1,2},...,d_{(k-1),r(k-1)},d_{k,rk},$ and $2m$ ancillary $q$-tokens $w_{1},w_{2},\cdots\!,w_{m},z_{1},...,z_{m}$.
@@
The labels on the directed arcs $x_{1},x_{2},\cdots,x_{4s-k}$ denote the variables that represent $q$-tokens on the incoming and outgoing arcs of transitions. The data $q$-tokens are assigned to places $P_{I1},P_{I2},...P_{Ik},$ and $P_{A}$ as in MISO quantum buffer, and not repeated here. When the input transition $T_{i}$ fires, $1\leq i\leq k,$ it forwards the current data and ancillary $q$-token, say $d_{i,y}$ and $w_{y}$ to place $P_{DA}$ i.e., $f_{1}$ becomes $d_{i,y},w_{y}$. When the output transition $T_{j}$ fires, $k+1\leq j\leq s,$ it forwards ancillary $q$-tokens $z_{y}$ and $w_{y}$ to $P_{A2},$ and data $q$-token $d_{i,y},$ to the corresponding output place $P_{oj}$. The behavior of the MIMO quantum buffer of capacity $m=2$ with two input and two output places $P_{I1},P_{I2},P_{O1}$ and $P_{O2}$ is illustrated in Figure 13. Input places $P_{I1}$ and $P_{I2}$ are initialized to 2 and 1 $q$-tokens, respectively.
Figure 13: The flow of $q$-tokens in a quantum MIMO buffer, $k=s=2,r_{1}=2,r_{2}=1,m=2.$
The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. We note that at the end of possible transitions, the three $q$-tokens may be distributed from input places $P_{I1},P_{I2},$ to output places $P_{O1}$ and $P_{O2}$ in one of six ways: $(P_{I1},P_{I2},P_{O1},P_{O2})=(0,1,2,0),(0,1,1,1),(0,1,0,2),(1,0,2,0),(1,0,1,1),(1,0,0,2)$ as the diagram in Figure 13 shows. Note that some transition sequences result in identical patterns of $q$-tokens in output places and are counted once. For example, both $T_{1},T_{3},T_{1},T_{4}$ and $T_{1},T_{4},T_{1},T_{3}$ result in the placement of one token in each of $P_{o1}$ and $P_{o1}.$ It takes $2m$ transitions, to process $m$ data $q$-tokens. Overall, the whole process is reversible, and can be repeated again.
@@ 4.5 Priority Quantum Buffer
A priority quantum buffer is a type of quantum buffer that prioritizes the processing of some $q$-tokens over others. The data $q$-tokens are divided into two groups $P_{I1},P_{I2},$ with those in $P_{I2}$ having a higher priority of transition over those in $P_{I1}.$ The priority is established using the ancillary $q$-tokens in $P_{A1}$ and $P_{A2}.$
@@
Figure 14 depicts the QPN model of a priority quantum buffer with capacity $m,$ four transitions $T_{1},T_{2},T_{3},T_{4}$, eight places $P_{I1},P_{I2},P_{A},P_{A1},P_{A2},P_{DA1},P_{DA2},P_{O}$, $r$ low priority data $q$-tokens, $d_{1},d_{2},...,d_{r}$, $n-r$ high priority data $q$-tokens $d_{r+1},d_{r+2},...,d_{n}$, $r+1\leq j\leq n,$ $m$ low priority ancillary $q$-tokens (qubits), $w_{1},w_{2},...,w_{m}$ and $m$ high priority ancillary $q$-tokens (qubits), $z_{1},z_{2},...,z_{m}$. The labels on the directed arcs $x_{1},x_{2},\cdots,x_{13}$ denote the variables that represent $q$-tokens on the incoming and outgoing arcs of transitions. It is assumed that the input transitions $T_{1},T_{2}$ can run independently. The directed arc $x_{8}$ serves as an inhibitor arc to prevent the firing of output transition $T_{3}$ when there is any $q$-token in $P_{DA2}$. The $q$-tokens are assigned to places $P_{I1},P_{I2},P_{A}$ and $P_{A1}$ and are initialized before the buffer starts operating as follows:
@@
Figure 15: The flow of $q$-tokens in a priority quantum buffer, $r=1,n=3,m=3.$
When the input transition, say $T_{i}$ fires, $1\leq i\leq 2,$ it forwards the current data and ancillary $q$-token, say $d_{y}$ and $z_{y}$ (or $w_{y}$) to place $P_{DAi}$ i.e., $f_{i}$ becomes either $d_{y},z_{y}$ (or $d_{y},w_{y}$). When the output transition, say $T_{j}$ fires, $3\leq j\leq 4,$ it forwards the corresponding ancillary $q$-token $z_{y}$ or $w_{y}$ to $P_{A2}$ and data $q$-token $d_{y}$ to $P_{O}$. However, the data $q$-tokens in place $P_{DA2}$ are prioritized over all tokens in place $P_{DA1}$ at the output place $P_{O}$ because transition $T_{3}$ cannot fire, due to the constraint of the inhibitor arc $x_{8}$. The behavior of the priority quantum buffer of capacity $m=4$ is illustrated in Figure 15. Input places $P_{I1}$ and $P_{I2}$ are initialized to one and two $q$-tokens, respectively. The quantities of $q$-tokens at the end of each transition can be verified by tracing the $q$-tokens as they move between places. We note that some transition sequences inhibit the firing of transition $T_{3}$ until after $T_{4}$ is fired. For example, when $T_{2}$ fires, it inhibits the firing of transition $T_{3}$ unless $T_{4}$ fired. It takes $2m$ transitions, to process $m$ data $q$-tokens. Overall, the whole process is reversible, and can be repeated again.
5 Quantum Runs and Simulations
In this section, we present the quantum circuit implementations of the two Q-SR flip-flops we described earlier on the IBM quantum machine (IBM Brisbane, QPU-Eagle-R3, Version 1.1.41) using IBM Composer [33]. We also validate the design of SISO, SIMO, MISO, MIMO, and priority quantum buffer on Jupyter [34] using Python and IBM Qiskit [32].